Tohban Report 2015-07-08

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 01 Jul 2015
End Date: 08 Jul 2015
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity remained relatively low this week although there were many C-class flares and 3 M-class flares later this week. Currently there are 9 ARs, and many of them are on the western side of the disk. We expect that solar activity remains relatively low next week.

How many GOES flares occurred?

 Flares above B, C, M, X class were      9    24     3     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      7    17     3     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      3     8     2     0

There were RHESSI flares/GOES flares 166 / 36 over the time range 01-Jul-15 07-Jul-15

Memory Management

The SSR fill level was not emptying at the end of daily pass sets. The daily average has gradually increased from 25% to 30% during this week.


Cold Plate 1 and 2 temperatures were 132.4 K and 131.0 K, respectively. Cold Tips 1 and 2 temperatures were 110.2 K and 108.2 K, respectively.

Spacecraft Status

Data Gaps

There was small (less than 900 sec) gaps. Some VC3 files were lost over the weekend.

Detector Issues

D4 front livetime has decreased quickly down to ~58% during the last few days, and so the front fast threshold was raised yesterday from 0x60 to 0x70.

D9 front livetime is even lower at ~52% but the rate of decrease is much slower. The fast threshold is unchanged.

D6 rear livetime has also decreased but gradually down to 30-40%. The rear fast threshold was raised yesterday from 0x68 to 0x70.

There were more light leaks in D2 front when compared to those of last week.

Attenuator Management

10 thin attenuator motions, 4 of them continued more than the minimum length. No thick attenuator motion.

There was a discussion about a possible experiment on attenuator behavior but it is likely to be carried out some weeks later.

Spacecraft Management

The decimation was changed from Active/Vigorous to Normal/Vigorous at 2015-07-01 19:34:56 UTC.

The front segment events during eclipses were turned off and took effect at 2015-07-01 18:53:45 UTC.

Decimation Normal/Vigorous
Night time data (fronts) No night events
Night time data (rears) No night events
Require extra passes? No
Attenuator operation Performing an experiment
Personal tools